Hi can you help me solve this equation please.
It says,
Solve the system by elimination. Show all your work for credit. You must show all your work.
{2x+y=25
{3y=2x-13
Thank you in advanced for any help!

Answer by ewatrrr(24785)

 
Hi,
2x+y=25
3y=2x-13
2x+y=25
-2x+3y= -13 |Adding 1st and 2nd EQ to eliminate x
4y = 12
y = 3 and x = 11 (2x+y=25, 2x + 3 = 25, 2x = 22, x = 11)
CHECKING our answer***
Using 3y=2x-13
9 = 22 - 13 = 9
Wish You the Best in your Studies.